russkaya05:

Simplify the following algebraic expression. 2(x - 9) + 6(-x + 2) + 4x anyone?

russkaya05:

am i correct

Hero:

You can tag people to your question to help.

russkaya05:

\[2(x-9) + 6(-x+2) + 4x\] Distribute: \[=(2)(x)+(2)(−9)+(6)(−x)+(6)(2)+4x\] \[=2x+−18+−6x+12+4x\] i combined Liked Terms: \[=2x+−18+−6x+12+4x\] \[=(2x+−6x+4x)+(−18+12)\] \[=−6\]

russkaya05:

?

Hero:

Correct

Hero:

Only thing I will say is to avoid writing + and - signs consecutively.

Hero:

Terms are separated by either a plus sign OR a minus sign

russkaya05:

k then thankyou

Hero:

yw